OpenStudy (anonymous):

What do I need to do to get \[\frac{ c }{ a }-\frac{ b^2 }{ 4a^2 }\]=\[\frac{ 4ac-b^2 }{ 4a }\]

OpenStudy (anonymous):

u cant lol

OpenStudy (anonymous):

these 2 expressions are not equal

OpenStudy (anonymous):

I mean equal, I am trying to prove the quadratic formula by completing the square.

OpenStudy (anonymous):

*dont

OpenStudy (anonymous):

and my text say the first expression can be manipulated into the second.

OpenStudy (anonymous):

\[\frac{ c }{ a }-\frac{ b^2 }{ 4a^2 }=\frac{4ac-b^2}{4a^2}\]

OpenStudy (anonymous):

ok how?

OpenStudy (anonymous):

@telltoamit

OpenStudy (anonymous):

c/a can be written as 4ac/4a^2

OpenStudy (anonymous):

how did this get a medal, I am know the wiser.

OpenStudy (anonymous):

hehee lol

hero (hero):

\[\frac{ c }{ a }-\frac{ b^2 }{ 4a^2 }\] 1. Multiply first fraction by 4a/4a to get: \[=\frac{4ac}{4a^2} -\frac{b^2}{4a^2}\] 2. Combine fractions to get: \[=\frac{4ac-b^2}{4a^2}\]3. You're finished